OpenStudy (anonymous):

Beginning at 6:50 a.m., planes leave an airport for New York every 45 minutes. You arrive at the airport at 10:15 a.m. How long will you have to wait for the next plane to New York? 15 minutes 20 minutes 25 minutes 30 minutes

OpenStudy (anonymous):

i was thinking d also

OpenStudy (dmezzullo):

I Meant To Say D Lol Typo.... :P

OpenStudy (anonymous):

7,17,27,37,47............197 means u get 20 numbers with 7 in between :)

OpenStudy (anonymous):

20

OpenStudy (amoodarya):

a(n)=7+10*n so solve for n a(n)<201 n is answer

OpenStudy (amoodarya):

sorry n+1 is answer because i start from 7

OpenStudy (texaschic101):

7,17,27,37,47,57,67,70,71,72,73,74,75,76,77,78,79,87,97,107,117,127,137,147,157,167,170,171,172,173,174,175,176,177,178,179,187,197 I count 40 7's......thats counting 77 and 177 as 4 sevens.

OpenStudy (anonymous):

um i got 20

OpenStudy (anonymous):

@texaschic101 welldone :0 but 72,74,76,78 arent odd numbers, if u can get them off, u will work the right answer :)

OpenStudy (texaschic101):

I must have read the question wrong....I thought it said all numbers with 7....srry ....the question is already gone and another one is up there.

OpenStudy (anonymous):

i forgot to include the 70,s odd numbers so now i have 28

OpenStudy (anonymous):

70's*

OpenStudy (texaschic101):

6:50 + :45 = 7:35 7:35 + :45 = 8:25 8:25 + :45 = 9:15 9:15 + :45 = 10:00 10:00 + :45 = 10:45 If you arrive at 10:15, the next plane leaves at 10:45....you will have to wait 30 minutes :)

OpenStudy (anonymous):

it was 20

OpenStudy (theviper):

First subtract :- 10:15-6:50=03:25hr {It is subtracted according to time} now change hrs into min 3:25hrs=3*60+25=180+25=205 Now divide it by 45 & the remainder will be answer :)

OpenStudy (texaschic101):

You might want to do it Vipers way, because the way I did it is too long....his is probably easier

OpenStudy (anonymous):

yup 20 cos the total time from 6:50 to 10:15 is 205 mints... and taking 6:50 start time the plane will be ready at 225 minuts cos that is the multiple of 45 the next departure time so u will have to wait 20 minutes

OpenStudy (texaschic101):

Wait...How is it not 30 minutes ???

OpenStudy (theviper):

no \(\Huge{\color{red}{25}}\) :)

OpenStudy (texaschic101):

Look at the way I did it.....did I just figure wrong ?

OpenStudy (theviper):

hey 7:35+:45=8:20 not 8:25 @texaschic101

OpenStudy (theviper):

Now no medal to me?? :D

OpenStudy (texaschic101):

ooohhhh....thank you for finding my mistake.....:)

OpenStudy (theviper):

LOL no problem :)

OpenStudy (texaschic101):

I will check my answers first before I post them...lol

OpenStudy (anonymous):

hey guys we r u stuck.. just count the time in minuts from 6:50 to 10:15..which is 205 minuts... and see the multiples of 45..nearest to 205 ..that is 225...so u have to wait 225-205=20 minute. its should be clear now if u tastse a biy for mathematics